# BRIGHT: A globally distributed multimodal building damage assessment dataset with very-high-resolution for all-weather disaster response

### v2

@@ Abstract
Disaster events occur around the world and cause significant damage to human life and property. Earth observation (EO) data enables rapid and comprehensive building damage assessment, an essential capability in the aftermath of a disaster to reduce human casualties and inform disaster relief efforts. Recent research focuses on the development of artificial intelligence (AI) models to achieve accurate mapping of unseen disaster events, mostly using optical EO data. However, solutions based on optical data are limited to clear skies and daylight hours, preventing a prompt response to disasters. Integrating multimodal EO data, particularly the combination of optical and synthetic aperture radar (SAR) imagery, makes it possible to provide all-weather, day-and-night disaster responses. Despite this potential, the development of robust multimodal AI models has been constrained by the lack of suitable benchmark datasets. In this paper, we present a Building damage assessment dataset using veRy-hIGH-resoluTion optical and SAR imagery (BRIGHT) to support AI-based all-weather disaster response. To the best of our knowledge, BRIGHT is the first open-access, globally distributed, event-diverse multimodal dataset specifically curated to support AI-based disaster response. It covers five types of natural disasters and two types of man-made disasters across 14 regions worldwide, with a particular focus on developing countries where external assistance is most needed. The optical and SAR imagery in BRIGHT, with a spatial resolution between 0.3 and 1 meters, provides detailed representations of individual buildings, making it ideal for precise damage assessment. In the experiments, we have tested seven advanced AI models trained with BRIGHT to validate the transferability and robustness. Beyond that, BRIGHT also serves as a challenging benchmark for a variety of tasks under disaster scenarios, including unsupervised domain adaptation (UDA), semi-supervised learning (SSL), unsupervised multimodal change detection (UMCD), and unsupervised multimodal image matching (UMIM). The experimental results serve as baselines to inspire future research and model development. The dataset, along with code and trained models, is available at https://github.com/ChenHongruixuan/BRIGHT and will be updated as new disaster data becomes available. BRIGHT also serves as the official dataset for the 2025 IEEE GRSS Data Fusion Contest Track II. We hope that this effort will promote the development of AI-driven EO methods in support of people in disaster-affected areas.
@@
A disaster is defined as a severe disruption in the functioning of a community or society due to the interaction between a hazard event and the conditions of exposure, vulnerability and capacity resulting in human, material, economic or environmental losses and impacts (Ge et al., 2020). According to the United Nations Office for Disaster Risk Reduction (UNDRR), between 1998 and 2017, natural disasters such as earthquakes, storms, and floods have affected 2.85 billion people and caused approximately 1.12 million deaths. These disasters have also resulted in economic losses of US$2,647 billion in disaster-affected countries (UNDRR, 2018a). The threat of disasters is likely to increase due to global urbanization (Kreibich et al., 2022; Bastos Moroz and Thieken, 2024). Rapid and comprehensive damage assessment is crucial in the aftermath of a disaster to make informed and effective rescue decisions to minimize losses and impacts. Building damage assessment aims to provide information, including the area and amount of damage, the rate of collapsed buildings, and the type of damage each building has incurred. This information is critical in the early stages of a disaster, as the distribution of damaged buildings is closely related to life-saving efforts in an emergency response (Xie et al., 2016; Adriano et al., 2021). Conducting field surveys after a disaster is difficult and dangerous, especially when transportation and communication systems are disrupted, preventing efficient on-site assessments. Earth observation (EO), also known as remote sensing, provides a safe and efficient way to obtain information on building damage in disaster areas because of its wide field of view and contactless operation.
@@
While accurate building damage maps can be obtained by human-expert visual interpretation of optical images, this process is time-consuming and labor-intensive for rapid large-scale assessments. In addition, it requires trained professionals. Therefore, recent literature has focused on developing automated methods for rapid building damage mapping (Tong et al., 2012; Xie et al., 2016; Gupta et al., 2019; Zheng et al., 2021). Among these, deep learning (DL) techniques have significantly improved the efficiency and accuracy of building damage assessment. Earlier work was toward a single disaster event with labels annotated for a particular disaster area to train a model, which is then used to generate building damage maps for the same event (Xie et al., 2016; Xia et al., 2023). However, since training data were limited to a few building types, damage patterns, and background land cover distributions, the resulting models mostly lack generalizability and struggle to produce accurate building damage maps for new disaster events, which limits their practical use. Recent large-scale benchmark datasets, for example, the xBD dataset (Gupta et al., 2019) of different types of disaster scenarios and damages, have made it possible to adopt DL models to quickly and accurately map building damages after a newly occurred, previously unseen disaster (Zheng et al., 2021; Shen et al., 2022; Guo et al., 2024; Wang et al., 2024; Chen et al., 2022a; Kaur et al., 2023; Chen et al., 2024). For example, Zheng et al. (2021) trained DL models on the xBD dataset and applied them to map the damages of buildings in two unseen human-made disaster events. These studies have demonstrated the effectiveness of DL models for building damage mapping.
@@
The optical EO technology uses a passive sensing technique, which requires solar illumination and cloud-free weather conditions. This severely limits the application of optical images in an emergency tool for all-weather disaster response (Adriano et al., 2021). In contrast, SAR sensors employ active illumination with longer microwaves and can acquire images in adverse weather conditions, offering great potential for all-weather disaster response. Most disaster events, especially wildfires, floods, and storms, are often accompanied by less-than-ideal imaging conditions. For example, Figure 1 shows EO imagery captured for a wildfire event that occurred in August 2023 in Hawaii, USA. The post-event optical image shown in Figure 1b does not provide clear surface information due to the effects of the wildfire smoke. However, the SAR image illustrated in Figure 1c is not affected by smoke and clearly shows the buildings damaged by the wildfire. Due to the advantages of SAR imagery, various SAR-based methods have been proposed for assessing building damage. These methods utilize intensity (Matsuoka and Yamazaki, 2005, 2010; Matsuoka et al., 2010), coherence (Yonezawa and Takeuchi, 2001; Arciniegas et al., 2007; Watanabe et al., 2016; Liu and Yamazaki, 2017), and polarization features (Yamaguchi, 2012; Chen and Sato, 2013; Watanabe et al., 2016; Karimzadeh and Mastuoka, 2017) to assess building damage at a block unit level, depending on the acquisition mode. Several studies have attempted to extend the block-level approach and have explored new approaches at the building instance level using higher spatial resolution sensors such as COSMO-SkyMed and TerraSAR-X (Liu et al., 2013; Brett and Guida, 2013; Chini et al., 2015; Ge et al., 2019). DL-based methods have also been explored with SAR data to assess building damage (Bai et al., 2018; Adriano et al., 2019; Bai et al., 2017; Li et al., 2023b). However, because of the lack of large-scale benchmark datasets, such as xBD in the optical domain, these methods have focused on local regions and single disaster events, and their ability to generalize to other disaster events remains largely unknown. Inherent challenges of SAR data, such as oblique viewing angles, speckle noise, object occlusion, and geometric distortions, complicate accurate building damage mapping compared to optical imagery (Brunner et al., 2010; Adriano et al., 2021). Furthermore, the limited availability of the VHR SAR data reduces its reliability as a source of pre-event data (Brunner et al., 2010). Considering these practical limitations, the most effective strategy for rapid assessment of building damage in all weather could arguably be to combine pre-event optical images, which provide accurate localization and detailed building information in the visible spectrum, with post-event SAR images, which capture structural information as a cue for building damage (Adriano et al., 2019). Previous methods have attempted to align the two modalities with traditional statistical models (Stramondo et al., 2006; Chini et al., 2009; Brunner et al., 2010; Wang and Jin, 2012). These statistical models are sensor-specific and require dedicated modeling for each sensor. DL methods offer a promising solution by automatically learning a high-dimensional feature space that aligns the two modalities. However, to train a DL model, one must have access to a high-quality large-scale dataset with comprehensive coverage of various disaster events and sufficient geographic diversity. This is a significant challenge that remains and needs to be addressed.
@@
To support DL research aimed at all-weather building damage mapping, we present BRIGHT, the first open and globally distributed multimodal VHR dataset for building damage assessment. Advances in EO technology have enabled data providers like Capella111https://www.capellaspace.com/ and Umbra222https://umbra.space/ to offer VHR SAR imagery at a sub-meter level resolution per pixel. This allows for detailed building damage assessments at the individual building level to guide targeted and effective rescue operations as required by emergency responders. Benefiting from the progress that has been made in EO, BRIGHT incorporates both pre-event optical imagery and post-event SAR imagery with spatial resolutions ranging from 0.3 meters to 1 meter per pixel. The types of disaster events considered in BRIGHT are earthquakes, storms (e.g., hurricane, cyclone), wildfires, floods, and volcanic eruptions. These natural disasters accounted for 84 of the fatalities and 94 of the economic losses between 1998 and 2017 (UNDRR, 2018a). In addition to natural disasters, the BRIGHT dataset further considers disasters caused by human activity, such as accidental explosions and armed conflicts, which also pose significant threats to human life and infrastructure and can occur unexpectedly, requiring a rapid response (UNDRR, 2018b). The events cover 14 different regions distributed around the globe with a focus on developing countries where external assistance is most urgently needed after a disaster. The labels are manually annotated with multi-level annotations that distinguish between damaged buildings and completely destroyed buildings.
@@ 0.1 Comparison with existing datasets
The comparison between BRIGHT and existing datasets for building damage assessment is summarized in Table 1. Most current building damage assessment datasets are limited in scale and scope due to limited available disaster events with corresponding open-source EO data and annotation effort (Rahnemoonfar et al., 2021; Gupta and Shah, 2021; Kaur et al., 2023). Because of the high cost and time required for pixel-level labeling, some of the existing datasets provide image-level labeling, indicating only whether an image contains damaged buildings (Fujita et al., 2017; Nguyen et al., 2017; Cheng et al., 2021; Xue et al., 2024). Although these image-level labeling datasets have served the community well, they lack the spatial precision needed to guide specific rescue operations. The xBD dataset (Gupta et al., 2019) is currently the largest open data collection, covering six natural disasters in 15 regions with more than 700,000 building instances. The xBD includes only optical EO data. Therefore, it does not support all-weather disaster response. Sun et al. (2024b) introduced a multimodal dataset, but it is limited to a single disaster event and contains only about 4,000 building instances. The small size makes it challenging to train DL models and limits the transferability of the trained models. The dataset most similar to BRIGHT is the BDD proposed by Adriano et al. (2021). The main differences between BDD and BRIGHT datasets are: (1) BRIGHT covers more disaster events and building instances, including both natural and human-made disasters. (2) BRIGHT has higher spatial resolution; while the highest resolution of SAR images in BDD is 1.2 meters, BRIGHT provides finer detail with spatial resolution from 0.3 meters to 1 meter to enable the detection of subtle structural damage in individual buildings. (3) Perhaps the most important difference is that whereas the re-distribution of BDD is restricted, BRIGHT is an open-source dataset publicly available to the global community. Apart from the datasets listed in Table 1, there are other datasets targeted at monitoring hazardous events related to disasters, including landslides (Ghorbanzadeh et al., 2022; Meena et al., 2023), floods (Bonafilia et al., 2020; Zhang et al., 2023) and wildfires (Artés et al., 2019; Huot et al., 2022; He et al., 2024), but are not concerned with building damage assessment.
@@ 0.2 Main contribution
We present BRIGHT, the first multimodal building damage dataset with a spatial resolution greater than 1 meter, which is publicly available to the community. BRIGHT employs a combination of pre-event optical imagery and post-event SAR imagery with various disaster events and rich geographic diversity to support the study of AI-based multimodal building damage mapping, especially in developing countries.
@@
We evaluate a suite of contemporary models on BRIGHT to establish robust baselines. Beyond supervised deep learning, BRIGHT can support a wide range of AI-based methods. It enables research in UDA, SSL, UMCD, and UMIM, among others. To demonstrate its utility, we benchmark a suite of representative models across several of these tasks. All experimental results, along with source code and pretrained model weights, are publicly released to provide strong baselines and accelerate future developments in EO-based disaster response.
@@ 1 Study areas and disaster events
We selected 14 disaster events across the globe for our BRIGHT dataset, as illustrated in Figure 2. Since both Capella and Umbra satellites were launched in 2020 and beyond, we focused on study areas where disasters occurred from 2020 onward. The selected regions are primarily in developing countries, where public administration and disaster response capacity tend to be weaker compared to developed nations, making international assistance more critical. The dataset covers five major types of natural disasters: earthquakes, storms (including hurricanes and cyclones), wildfires, floods, and volcanic eruptions. Additionally, it includes human-made disasters, such as accidental explosions and armed conflicts.
| Disaster area | Type of disaster | Date | GSD (m/pixel) | Provider | No. of tiles | No. of building |
@@ 1.1 Explosion in Beirut, 2020
On August 4, 2020, a massive explosion occurred at the Port of Beirut in Lebanon, caused by the improper storage of 2,750 tons of ammonium nitrate. The explosion caused widespread damage within a radius of several kilometers, significantly impacting the port and the surrounding neighborhoods, including areas such as Gemmayzeh, Mar Mikhael, and Achrafieh. It resulted in 218 deaths, more than 7,000 injuries, and left approximately 300,000 people homeless. Economic losses were estimated to be between 10 billion and 15 billion USD. The disaster compounded Lebanon’s ongoing economic challenges and contributed to political instability and social unrest.

### v3

@@ Abstract
Disaster events occur around the world and cause significant damage to human life and property. Earth observation (EO) data enables rapid and comprehensive building damage assessment, an essential capability in the aftermath of a disaster to reduce human casualties and inform disaster relief efforts. Recent research focuses on the development of artificial intelligence (AI) models to achieve accurate mapping of unseen disaster events, mostly using optical EO data. However, solutions based on optical data are limited to clear skies and daylight hours, preventing a prompt response to disasters. Integrating multimodal EO data, particularly the combination of optical and synthetic aperture radar (SAR) imagery, makes it possible to provide all-weather, day-and-night disaster responses. Despite this potential, the development of robust multimodal AI models has been constrained by the lack of suitable benchmark datasets. In this paper, we present a Building damage assessment dataset using veRy-hIGH-resoluTion optical and SAR imagery (Bright) to support AI-based all-weather disaster response. To the best of our knowledge, Bright is the first open-access, globally distributed, event-diverse multimodal dataset specifically curated to support AI-based disaster response. It covers five types of natural disasters and two types of human-made disasters across 14 regions worldwide, with a particular focus on developing countries where external assistance is most needed. The optical and SAR imagery in this dataset, with a spatial resolution between 0.3 and 1 meters, provides detailed representations of individual buildings, making it ideal for precise damage assessment. In the experiments, we have tested seven advanced AI models trained with Bright to validate the transferability and robustness. Beyond that, it also serves as a challenging benchmark for a variety of tasks under disaster scenarios, including unsupervised domain adaptation (UDA), semi-supervised learning (SSL), unsupervised multimodal change detection (UMCD), and unsupervised multimodal image matching (UMIM). The experimental results serve as baselines to inspire future research and model development. The dataset, along with code and trained models, is available at https://github.com/ChenHongruixuan/BRIGHT and will be updated as new disaster data becomes available. Bright also serves as the official dataset for the 2025 IEEE GRSS Data Fusion Contest Track II. We hope that this effort will promote the development of AI-driven methods in support of people in disaster-affected areas.
@@
A disaster is defined as a severe disruption in the functioning of a community or society due to the interaction between a hazard event and the conditions of exposure, vulnerability and capacity resulting in human, material, economic or environmental losses and impacts (Ge et al., 2020). According to the United Nations Office for Disaster Risk Reduction (UNDRR), between 1998 and 2017, natural disasters such as earthquakes, storms, and floods have affected 4.4 billion people and caused approximately 1.3 million deaths. These disasters have also resulted in economic losses of US$2,647 billion in disaster-affected countries (UNDRR, 2018a). The threat of disasters is likely to increase due to global urbanization (Kreibich et al., 2022; Bastos Moroz and Thieken, 2024). Rapid and comprehensive damage assessment is crucial in the aftermath of a disaster to make informed and effective rescue decisions to minimize losses and impacts. Building damage assessment aims to provide information, including the area and amount of damage, the rate of collapsed buildings, and the type of damage each building has incurred. This information is critical in the early stages of a disaster, as the distribution of damaged buildings is closely related to life-saving efforts in an emergency response (Xie et al., 2016; Adriano et al., 2021). Conducting field surveys after a disaster is difficult and dangerous, especially when transportation and communication systems are disrupted, preventing efficient on-site assessments. Earth observation (EO), also known as remote sensing, provides a safe and efficient way to obtain information on building damage in disaster areas because of its wide field of view and contactless operation.
@@
While accurate building damage maps can be obtained by human-expert visual interpretation of optical images, this process is time-consuming and labor-intensive for rapid large-scale assessments. In addition, it requires trained professionals. Therefore, recent literature has focused on developing automated methods for rapid building damage mapping (Tong et al., 2012; Xie et al., 2016; Gupta et al., 2019; Zheng et al., 2021). Among these, machine learning (ML) and deep learning (DL) techniques have significantly improved the efficiency and accuracy of building damage assessment. Earlier work was toward a single disaster event with labels annotated for a particular disaster area to train a model, which is then used to generate building damage maps for the same event (Xie et al., 2016; Xia et al., 2023). However, since training data were limited to a few building types, damage patterns, and background land cover distributions, the resulting models mostly lack generalizability and struggle to produce accurate building damage maps for new disaster events, which limits their practical use. Recent large-scale benchmark datasets, for example, the xBD dataset (Gupta et al., 2019) of different types of disaster scenarios and damages, have made it possible to adopt DL models to quickly and accurately map building damages after a newly occurred, previously unseen disaster (Zheng et al., 2021; Chen et al., 2022a; Shen et al., 2022; Kaur et al., 2023; Guo et al., 2024; Wang et al., 2024; Chen et al., 2024). For example, Zheng et al. (2021) trained DL models on the xBD dataset and applied them to map the damages of buildings in two unseen human-made disaster events. These studies have demonstrated the effectiveness of DL models for building damage mapping.
@@
The optical EO technology uses a passive sensing technique, which requires solar illumination and cloud-free weather conditions. This severely limits the application of optical images in an emergency tool for all-weather disaster response (Adriano et al., 2021). In contrast, SAR sensors employ active illumination with longer microwaves and can acquire images in adverse weather conditions, offering great potential for all-weather disaster response. Most disaster events, especially wildfires, floods, and storms, are often accompanied by less-than-ideal imaging conditions. For example, Figure 1 shows EO imagery captured for a wildfire event that occurred in August 2023 in Hawaii, USA. The post-event optical image shown in Figure 1b does not provide clear surface information due to the effects of the wildfire smoke. However, the SAR image illustrated in Figure 1c is not affected by smoke and clearly shows the buildings damaged by the wildfire. Due to the advantages of SAR imagery, various SAR-based methods have been proposed for assessing building damage. These methods utilize intensity (Matsuoka and Yamazaki, 2005, 2010; Matsuoka et al., 2010), coherence (Yonezawa and Takeuchi, 2001; Arciniegas et al., 2007; Watanabe et al., 2016; Liu and Yamazaki, 2017), and polarization features (Yamaguchi, 2012; Chen and Sato, 2013; Watanabe et al., 2016; Karimzadeh and Mastuoka, 2017) to assess building damage at a block unit level, depending on the acquisition mode. Several studies have attempted to extend the block-level approach and have explored new approaches at the building instance level using higher spatial resolution sensors such as COSMO-SkyMed and TerraSAR-X (Liu et al., 2013; Brett and Guida, 2013; Chini et al., 2015; Ge et al., 2019). DL-based methods have also been explored with SAR data to assess building damage (Bai et al., 2018; Adriano et al., 2019; Bai et al., 2017; Li et al., 2023b). However, because of the lack of large-scale benchmark datasets, such as xBD in the optical domain, these methods have focused on local regions and single disaster events, and their ability to generalize to other disaster events remains largely unknown. Inherent challenges of SAR data, such as oblique viewing angles, speckle noise, object occlusion, and geometric distortions, complicate accurate building damage mapping compared to optical imagery (Adriano et al., 2021; Xia et al., 2025). Furthermore, the limited availability of the VHR SAR data reduces its reliability as a source of pre-event data (Brunner et al., 2010; Adriano et al., 2021). Considering these practical limitations, the most effective strategy for rapid assessment of building damage in all weather could arguably be to combine pre-event optical images, which provide accurate localization and detailed building information in the visible spectrum, with post-event SAR images, which capture structural information as a cue for building damage (Adriano et al., 2019). Previous methods have attempted to align the two modalities with traditional statistical models (Stramondo et al., 2006; Chini et al., 2009; Brunner et al., 2010; Wang and Jin, 2012). These statistical models are sensor-specific and require dedicated modeling for each sensor. DL methods offer a promising solution by automatically learning a high-dimensional feature space that aligns the two modalities. However, to train a DL model, one must have access to a high-quality large-scale dataset with comprehensive coverage of various disaster events and sufficient geographic diversity. This is a significant challenge that remains and needs to be addressed.
@@
To support DL research aimed at all-weather building damage mapping, we present Bright, the first open and globally distributed multimodal VHR dataset for building damage assessment. Advances in EO technology have enabled data providers like Capella111https://www.capellaspace.com/ and Umbra222https://umbra.space/ to offer VHR SAR imagery at a sub-meter level resolution per pixel. This allows for detailed building damage assessments at the individual building level to guide targeted and effective rescue operations as required by emergency responders. Benefiting from the progress that has been made in EO, Bright incorporates both pre-event optical imagery and post-event SAR imagery with spatial resolutions ranging from 0.3 meters to 1 meter per pixel. The types of disaster events considered in Bright are earthquakes, storms (e.g., hurricane, cyclone), wildfires, floods, and volcanic eruptions. These natural disasters accounted for 84 of the fatalities and 94 of the economic losses between 1998 and 2017 (UNDRR, 2018a). In addition to natural disasters, the Bright dataset further considers disasters caused by human activity, such as accidental explosions and armed conflicts, which also pose significant threats to human life and infrastructure and can occur unexpectedly, requiring a rapid response (UNDRR, 2018b; Dietrich et al., 2025). The 14 disaster events cover 23 different regions distributed around the globe with a focus on developing countries where external assistance is most urgently needed after a disaster. The labels are manually annotated with multi-level annotations that distinguish between damaged buildings and completely destroyed buildings.
@@ 0.1 Comparison with existing datasets
The comparison between Bright and existing datasets for building damage assessment is summarized in Table 1. Most current building damage assessment datasets are limited in scale and scope due to limited available disaster events with corresponding open-source EO data and annotation effort (Rahnemoonfar et al., 2021; Gupta and Shah, 2021; Kaur et al., 2023). Because of the high cost and time required for pixel-level labeling, some of the existing datasets provide image-level labeling, indicating only whether an image contains damaged buildings (Fujita et al., 2017; Nguyen et al., 2017; Cheng et al., 2021; Xue et al., 2024). Although these image-level labeling datasets have served the community well, they lack the spatial precision needed to guide specific rescue operations. The xBD dataset (Gupta et al., 2019) is currently the largest open data collection, covering six natural disasters in 15 regions with more than 700,000 building instances. However, the xBD includes only optical EO data. It does not support all-weather disaster response. Sun et al. (2024b) introduced a multimodal dataset, but it is limited to a single disaster event and contains only about 4,000 building instances. The small size makes it challenging to train DL models and limits the transferability of the trained models. The dataset most similar to Bright is the BDD proposed by Adriano et al. (2021). The main differences between BDD and Bright datasets are: (1) Bright covers more disaster events and building instances, including both natural and human-made disasters. (2) Bright has higher spatial resolution; while the highest resolution of SAR images in BDD is 1.2 meters, Bright provides finer detail with spatial resolution from 0.3 meters to 1 meter to enable the detection of subtle structural damage in individual buildings. (3) Perhaps the most important difference is that whereas the re-distribution of BDD is restricted, Bright is an open-source dataset publicly available to the global community. Apart from the datasets listed in Table 1, there are other datasets targeted at monitoring hazardous events related to disasters, including landslides (Ghorbanzadeh et al., 2022; Meena et al., 2023), floods (Bonafilia et al., 2020; Zhang et al., 2023) and wildfires (Artés et al., 2019; Huot et al., 2022; He et al., 2024), but are not concerned with building damage assessment.
@@ 0.2 Main contribution
We present Bright, the first multimodal building damage dataset with sub-meter spatial resolution, which is publicly available to the community. Bright employs a combination of pre-event optical imagery and post-event SAR imagery with various disaster events and rich geographic diversity to support the study of AI-based multimodal building damage mapping, especially in developing countries.
@@
We evaluate a suite of contemporary models on Bright to establish robust baselines. Beyond supervised deep learning, Bright can support a wide range of AI-based methods. It enables research in UDA, SSL, UMCD, and UMIM, among others. To demonstrate its utility, we benchmark a suite of representative models across several of these tasks. All experimental results, along with source code and pretrained model weights, are publicly released to provide strong baselines and accelerate future developments for disaster response in the community.
@@ 1 Study areas and disaster events
We selected 14 disaster events across the globe for Bright, as illustrated in Figure 2. Since both Capella and Umbra satellites were launched in 2020 and beyond, we focused on study areas where disasters occurred from 2020 onward. The selected regions are primarily in developing countries, where public administration and disaster response capacity tend to be weaker compared to developed nations, making international assistance more critical. The dataset covers five major types of natural disasters: earthquakes, storms (including hurricanes and cyclones), wildfires, floods, and volcanic eruptions. Additionally, it includes human-made disasters, such as accidental explosions and armed conflicts.
| Disaster area | Type of disaster | Date | GSD (m/pixel) | Data provider / source | No. of tiles | No. of building |
@@ 1.1 Explosion in Beirut, 2020
On August 4, 2020, a massive explosion occurred at the Port of Beirut in Lebanon, caused by the improper storage of 2,750 tons of ammonium nitrate. The explosion caused widespread damage within a radius of several kilometers, significantly impacting the port and the surrounding neighborhoods, including areas such as Gemmayzeh, Mar Mikhael, and Achrafieh. It resulted in 218 deaths, more than 7,000 injuries, and left approximately 300,000 people homeless (Fakih and Majzoub, 2021). Economic losses were estimated to be 15 billion USD (Valsamos et al., 2021). The disaster compounded Lebanon’s ongoing economic challenges and contributed to political instability and social unrest.